### v1

# A Time Series Analysis of Assertions in the Linux Kernel

Abstract.
Assertions are a classical and typical software development technique. These are extensively used also in operating systems and their kernels, including the Linux kernel. The paper patches a gap in existing knowledge by empirically examining the longitudinal evolution of assertion use in the Linux kernel. According to the results, the use of assertions that cause a kernel panic has slightly but not substantially decreased from the kernel’s third to the sixth release series. At the same time, however, the use of softer assertion variants has increased; these do not cause a panic by default but instead produce warnings. With these time series results, the paper contributes to the existing but limited empirical knowledge base about operating system kernels and their long-term evolution.
Design by contract; kernel panic; software evolution; subsystems
ccs: Software and its engineering Operating systems
1. Introduction
Assertions are a classical programming technique used to validate a program’s execution at runtime. They also serve to document code for other developers. Assertions are typically used within functions to test preconditions placed at the beginning of a function and postconditions placed at the end of the function, although constraints can be placed also upon return values and intermediate states of the function (Rosenblum95, ). In software engineering and the computing world in general, assertions are also behind a classical debate between so-called defensive programming and design by contract.
In general, the defensive variant seeks to particularly validate each and every input, whereas the latter assumes an implicit contract made explicit through assertions between a caller and a supplier. In terms of the design by contract paradigm, a supplier can assume that certain conditions are always guaranteed by a caller, whereas the caller can assume a safe output provided by the supplier. When operating with assertions instead of defensive programming practices, “any runtime violation of an assertion is not a special case but always the manifestation of a software bug”, tracing of which can be started by examining the given preconditions, implying “a bug in the client (caller)”, and the given postconditions associated with “a bug in the supplier (routine)” (Meyer92, , p. 44). Regardless whichever side one takes in this debate, assertions are frequently used—and triggered—also in operating systems and their kernels. This background motives the paper’s focus on the Linux kernel.
With respect to the Linux kernel, assertions have frequently been observed to be behind many of the crashes triggered by continuous kernel fuzzing frameworks (Ruohonen19RSDA, ). In the kernel context an explicitly triggered crash corresponds with a kernel panic. Without an assertion causing a panic, an alternative behavior might be that a whole operating system just hangs. Also these cases have historically been common in the Linux kernel (Gu03, ; Yoshimura12, ). Thus, the design by contract paradigm has a slightly different interpretation in the kernel context within which failing spectacularly may often, or at least sometimes, be a better option than to continue execution in undefined waters (FreeBSD24, ). Different fallback and recovery solutions are often also difficult to design and implement in the kernel context.
These and related points motivated a lively debate recently among kernel developers and other open source developers. The motivating argument put forward was that Linux kernel developers should slowly move away from using assertions that cause an outright panic; instead, they should prefer warnings and try to come up with means to recover from fatal errors (LWN24a, ). In the Linux kernel the former panic-inducing assertions are used with different BUG()-style macros, whereas the latter are typically used with different WARN()-style variants. For a lack of a better terminology, these are hereafter known as “hard” and “soft” assertions. Regarding the hard assertions, the kernel documentation argues that “‘I’m too lazy to do error handling’ is not an excuse for using BUG()”, and that major “internal corruptions with no way of continuing may still use BUG(), but need good justification” (Kernel24a, ). With this guidance and the debate in mind, two research questions (RQs) are examined:
  • RQ.1: Has the kernel’s growing size been met with a growing use of soft assertions in some or all of its notable subsystems?
  • RQ.2: Has the amount of hard assertion declined over time in some or all of the notable subsystems of the Linux kernel?
Regarding the paper’s structure, there are four sections remaining. The first of these, Section 2, briefly discusses related work from a broader perspective. Then, Section 3 elaborates the research setup, including the empirical data used and the methods for its analysis. Results are presented in Section 4. The final Section 5 concludes.
2. Related Work
In addition to the works already referenced, a couple of larger knowledge bases are worth mentioning. These allow to also better motivate the RQs. Thus, regarding RQ.1, the software evolution research domain provides a good reference point. According to the so-called laws of software evolution, which were developed by Lehman and associates from the 1970s onward (Lehman97, ), many—if not the most—software artifacts are continuously changing, continuously growing, and continuously getting more complex (Ruohonen15JSEP, ). These laws correlate with the more recent concept of technical debt; it is necessary to pay back maintenance debt in order to continue to have a given software artifact in a reasonable state in the face of continuous growth and complexity. The Linux kernel has been seen as a prime example supporting the Lehman’s laws (Israeli10, ). Regarding growth, particularly new device drivers continuously added have been the driving force (Chou01, ; Ruohonen24EASE, ). Partially due to the growth, the Linux kernel is also continuously and heavily refactored, as manifested by the extensive amount of code churn (Ruohonen19RSDA, ). In other words, technical debt is also continuously paid back. These points provide a motivation also for RQ.1. Although developers are encouraged to move away from hard assertions, retaining the existing but perhaps implicit design by contract principle requires that also soft assertions should grow alongside the growth of the kernel’s code base.
The second knowledge base is the somewhat limited amount of existing empirical research on assertions. This claim should not be interpreted to mean that empirical research would be absent; particularly the benefits of assertions in software testing have been investigated empirically (Taromirad23, ). In this regard, assertions have been observed to correlate with the effectiveness of test suites (Zhang15, ) and less buggy files (Kudrjavets06, ), among other things. Another empirical research area has focused on the benefits of assertions for troubleshooting and debugging, understandability, and documentation (Patel17, ; Peruma24, ). However, to the best of the author’s knowledge, no empirical studies have been conducted about the longitudinal evolution of the use of assertion in the Linux kernel or any other kernel for that matter. Thus, the paper fills a small but notable gap in existing knowledge.
3. Research Setup
3.1. Data and Time Series
The Linux kernel more or less follows the traditional semantic versioning scheme. Release series are the highest level of abstraction; currently, the Linux kernel operates with the sixth release series. Major releases, such as $6.1,6.2,6.3,\ldots$, are made based on a loosely time-based release process; a new major release occurs roughly every two or three months (Kernel24b, ). The adjective major signifies that such releases may contain everything from new features to internal changes in the kernel’s application programming interface. Each major release typically changes several hundred thousand lines of code (LOC). In addition, each major release is maintained by a separate team who manages the release engineering of minor releases, such as $6.1.1,6.1.2,6.1.3$, and so forth. These minor releases typically contain only bug fixes. Therefore, the major releases are more suitable for observing the evolution of assertions in the kernel.
Thus, each major release in the third, fourth, fifth, and sixth release series was downloaded from the kernel’s online archives (Kernel24c, ). At the time of writing (20 December 2024), the amounts of major releases released in these release series were 20, 21, 20, and 13, respectively. Thus, the length of all time series constructed is 74 releases. Although the length is rather short in general, it is still sufficient for time series analysis with formal methodologies.
Both the overall software architecture of the Linux kernel and its software development rely on different subsystems (Ruohonen24EASE, ). In what follows, the time series analysis is also based on major subsystems of the kernel. The major ones observed are: arch (architectural code dependent on instruction set architectures), drivers (device drivers), fs (file systems), net (networking), and sound (sound architectures). By following existing research (JiangJiang24, ; Ruohonen24EASE, ), these were proxied by the top-level directories in the kernel’s file system representation. In addition, a category of others was used to merge other notable subsystems, such as kernel for general kernel code, crypto for cryptographic implementations, and security for the kernel’s security implementations. In addition, this leftover category contains all other top-level directories insofar as they contained C programming code, as identified by files ending to a .c suffix.
The reason for the leftover category (or, hereafter, a “subsystem” in order to simplify the language) is that the kernel’s subsystems vary considerably in terms of their sizes. This point is evident from Table 1, which shows the average sizes of the subsystems in terms of .c-files contained in these and the total amount of LOC, including comments, in these files, scaled by the seventy-four releases sampled. Even with the merging of top-level directories, others is smaller than the other subsystems in terms of LOC. The device driver subsystem is the largest one regardless of a size metric.
Table 1. Average Sizes of the Subsystems Across Releases
Subsystem Files LOC
arch $4,834$ $1,590,037$
drivers $14,432$ $11,692,942$
fs $1,224$ $1,146,606$
net $1,321$ $964,467$
sound $1,253$ $931,181$
others $2,438$ $82,114$
Then, for each subsystems, four time series were constructed:
  1. (1)
    $\textmd{FILE}_{it}$ counts the number of files containing C programming code, as identified by a file’s ending to a suffix .c, in the $i$:th subsystem of the $t$:th major release.
  2. (2)
    $\textmd{LOC}_{it}$ measures the lines of code, including code comments, of the files containing C code, as identified similarly to $\textmd{FILES}_{it}$, in the $i$:th subsystem of the $t$:th major release.
  3. (3)
    $\textmd{BUG}_{it}$ measures the amount of hard assertions in the C files present in the $i$:th subsystem of the $t$:th major release. These were identified by parsing each file line by line and counting the occurrences of the character strings BUG( and BUG_ON(.
  4. (4)
    $\textmd{WARN}_{it}$ measures the amount of soft assertions in the .c-suffixed files present in the $i$:th subsystem of the $t$:th major release. Analogously to $\textmd{BUG}_{it}$, these were identified by parsing each file line by line and counting the occurrences of the character strings WARN(, WARN_ON(, and WARN_ON_ONCE(.
The computations are carried out with the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series as well scaled variants of the two time series. Two simple scaling functions are used: the first scales the plain series by the total amount of hard and soft assertions in the $i$:th subsystems of the $t$:th release, and the second by a subsystem’s size at a given major release. In other words:
(1) $\overline{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{BUG}_{it}+\textmd{WARN}_{it}}\quad\textmd{and}\quad\widetilde{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{FILE}_{it}},$
and likewise for $\textmd{WARN}_{it}$. The scaling by $\textmd{FILE}_{it}$ is arbitrary; none of the results reported changed when scaled by $\textmd{LOC}_{it}$. Finally, it can be remarked that a replication package is available online (Ruohonen24rep, ).
3.2. Methods
Statistical methods for cumulative sum (CUSUM) processes are suitable for examining the research questions. The history behind CUSUM processes is rich, starting from the 1950s and peaking in the 1970s during which first formal statistical tests were derived for testing structural changes in time series (Brown75, ). By a structural change it is meant that a model estimated for a time series does not remain stable over time. The concept is much broader than the concept of (covariance) stationary, which only requires that a mean and a variance of a time series remain constant. However, in the present context these two concepts are closely related because the focus is on the means of $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. With respect to RQ.2, a weak requirement is that the mean of the former should not remain constant. A strong requirement implies that the means of both series should not remain constant because developers are encouraged to move from hard assertions to the softer variants.
The CUSUM processes are thus computed for both “static mean models” and “dynamic autoregression models”, both estimated with standard ordinary least squares (OLS). To clarify these modeling concepts, the former models merely estimate $\textmd{BUG}_{it}=\alpha_{t}+\epsilon_{t}$ and $\textmd{WARN}_{it}=\beta_{t}+\varepsilon_{t}$, where $\alpha_{t}$ and $\beta_{t}$ are intercepts, and the residuals $\epsilon_{t}$ and $\varepsilon_{t}$ are assumed to have zero means and variances $\sigma^{2}_{\epsilon_{t}}$ and $\sigma^{2}_{\varepsilon_{t}}$ for $t=1,\ldots,74$. The underlying (null) hypotheses are that $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{74}$ and $\sigma^{2}_{\epsilon_{1}}=\sigma^{2}_{\epsilon_{2}}=\cdots=\sigma^{2}_{\epsilon_{74}}$, and likewise for $\beta_{t}$ and $\sigma^{2}_{\varepsilon_{t}}$. In particular, the conditional means should remain constant, although the word conditional is a slight misnomer in the present context because only constants are included in the models.
The dynamic first-order autoregression, denoted by AR(1), models are estimated for differenced series, $\Delta\textmd{BUG}_{it}=\textmd{BUG}_{it}-\textmd{BUG}_{it-1}$ and $\Delta\textmd{WARN}_{it}$, given that many of the series are not stationary. For each subsystem $i=1,\ldots,6$, a simple autoregression model $\Delta\textmd{BUG}_{it}=\phi_{t}+\rho_{t}\Delta\textmd{BUG}_{it-1}+u_{t}$ is estimated, and likewise for $\Delta\textmd{WARN}_{it}$. The underlying hypothesis again is that $\phi_{t}$, $\rho_{t}$, and the variance $\sigma^{2}_{u_{t}}$ of the residual term $u_{t}$ remain constant throughout the seventy-four major releases of the Linux kernel. The same applies to the corresponding parameters in the AR(1) model for $\Delta\textmd{WARN}_{it}$.
The CUSUM processes are particularly useful as visual diagnostic tools. The strucchange package (Zeileis02, ) for R is used for the visualizations and associated computations. It allows to visualize the CUSUM processes computed with visual boundaries; 99% confidence level boundaries are used for the visualizations presented. In addition, the package offers more formal so-called M-fluctuation tests (Zeileis06, ). These are used to compute analogous models than the static (OLS) mean models used for the visualizations except that a Poisson regression is used for the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series. The four scaled series $\overline{\textmd{BUG}}_{it}$, $\widetilde{\textmd{BUG}}_{it}$, $\overline{\textmd{WARN}}_{it}$, and $\widetilde{\textmd{WARN}}_{it}$ are tested with OLS because they do not represent count data any more.
In addition, to examine RQ.1 in a little more detail, a simple non-linear growth curve estimation is carried out. Although numerous growth curves have been specified in the literature (Ruohonen15COSE, ), the standard logistic growth curve is sufficient for the present purposes. It is written the growthcurver package (Sprouffske16, ) used for computation as:
(2) $y_{t}=\frac{\alpha}{1+\frac{\alpha-y_{0}}{y_{0}}e^{-\beta t}},$
where $y_{t}$ denotes a time series estimated with non-linear least squares, $\alpha$ is the asymptotic maximum, $y_{0}$ is the initial value for a growth, and $\beta$ is the growth rate. The interest is not to examine the statistical performance of the logistic growth curves estimated, but rather to examine the growth rate parameters across the time series. With respect to RQ.1, an estimated $\hat{\beta}$ for $\textmd{WARN}_{it}$ should be roughly comparable to equivalent parameters estimated for $\textmd{FILE}_{it}$ and $\textmd{LOC}_{it}$ in all $i=[1,6]$ subsystems. In this case the kernel’s pace of growth would be matched by use of soft assertions at a roughly same speed. It should be mentioned that some of the series are not increasing but decreasing—as is also presumed by RQ.2. As the logistic growth curve is ill-suited for such series, the growth rate coefficients are interpreted together with their standard errors.
4. Results
The presentation of the results can be started by taking a look at the time series across the six subsystems. Three main series are thus shown in Fig. 1. The conventional wisdom holds well: the drivers subsystem has grown considerably in terms of new C programming files added throughout the releases. Although the magnitudes are much smaller, also the sound and others subsystems have witnessed accelerating growth. Also the fs and net subsystems have grown, although at a much slower pace and smaller magnitudes. The arch subsystem is the only one in which the amount of .c-suffixed files has decreased. Although a reason can be left for further examinations, the observation could be interpreted in a positive light in a sense that machine-dependent code has slightly decreased.
Figure 1. $\textmd{FILE}_{it}$, $\textmd{BUG}_{it}$, and $\textmd{WARN}_{it}$
Table 2. Logistic Growth Rate Parameters and Their Standard Errors (SEs)1
$\textmd{FILE}_{it}$ $\textmd{LOC}_{it}$ $\textmd{BUG}_{it}$ $\textmd{WARN}_{it}$
$\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$)
Subsystem arch $<$0.0001 $<$0.0001 0.0534 0.0015
drivers 0.0739 0.0028 0.0648 0.0026 0.0676 0.0027
fs 0.0306 0.0033 0.0413 0.0028 $<$0.0001 $<$0.0001 0.0553 0.0033
net 0.0781 0.0044 0.0641 0.0024 $<$0.0001 $<$0.0001 0.0606 0.0020
sound 0.0574 0.0028 0.0594 0.0022 0.2722 0.0336
others 0.0649 0.0011 0.0588 0.0016 0.0677 0.0060 0.0580 0.0018
  • 1 Estimates are not shown in case a standard error is larger than a parameter estimate.
Table 3. Generalized M-fluctuation Tests ($p$-values)
Series $\textmd{BUG}_{it}$ $\overline{\textmd{BUG}}_{it}$ $\widetilde{\textmd{BUG}}_{it}$ $\textmd{WARN}_{it}$ $\overline{\textmd{WARN}}_{it}$ $\widetilde{\textmd{WARN}}_{it}$
Model Poisson OLS OLS Poisson OLS OLS
Subsystem arch $0.0003$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
drivers $0.0002$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
fs $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
net $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
sound $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
others $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $0.0001$
Figure 2. OLS-CUSUM Processes for $\textmd{FILE}_{it}$ and $\textmd{BUG}_{it}$ (the circled boundaries denote a $p<0.001$ confidence level)
Figure 3. OLS-CUSUM Processes for AR(1) models of $\Delta\textmd{FILE}_{it}$ and $\Delta\textmd{BUG}_{it}$ (the boundaries denote a $p<0.001$ confidence level)
With these observations about growing sizes of most subsystems, the logistic growth curve estimates can be subsequently considered. The growth rate parameters estimated are thus shown in Table 2. As was noted in the previous Section 3.2, some of the estimates are not reasonable, but regarding those that are, it is important to emphasize that the coefficients for $\textmd{WARN}_{it}$ are in the majority of subsystems well in line with those estimated for the two time series $\textmd{FILE}_{it}$ and $\textmd{LOC}_{t}$ measuring software size. In fact, the growth rates of soft assertion use have in some subsystems surpassed the growth rates of software size. Furthermore, in the fs subsystem and particularly in the net subsystem the ratio $\textmd{WARN}_{it}~{}/~{}\textmd{FILE}_{it}$ has exceeded a value one. In general, the answer to RQ.1 is positive; in most major subsystems the growing sizes have been met with equally growing uses of soft assertions. Having said that, it remains a subject of a debate whether the growth rates of $\textmd{WARN}_{it}$ should be even higher in order to satisfy the ever-growing sizes of some subsystems, the implicit design by contract principle, and the overall goal of providing non-fatal debugging information to users.
Turning to the CUSUM processes, the formal M-fluctuation tests are summarized in Table 3. It can be concluded that all static mean models reject the null hypotheses of stable processes, no matter of a particular statistical significance level specified or preferred. When recalling the negation discussed in Section 3.2, the result is a positive finding with respect to both RQ.1 and RQ.2. In other words, neither the use of hard assertions nor the use of soft assertions has remained stable over the years. However, taking a visual look at the underlying CUSUM processes reveals two exceptions.
As can be seen from Fig. 3, the static mean models for the $\textmd{BUG}_{it}$ time series mostly stay within the circled boundaries in the arch and drivers subsystems. Only minor boundary crossings are present for these two time series. In all other subsystems the boundaries are heavily crossed by the static mean models for both $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. Although there is no right or wrong answer to a question whether one should prefer the visual inspection over the formal M-fluctuation tests, it can be concluded that the answer to RQ.2 is not entirely uniform. Given the question’s phrasing “some or all”, the conclusion is on the side of the former word. This conclusion warrants a brief speculation on what might explain the stable use of hard assertions in the two outlying subsystems across time. A reasonable presumption might be that both arch and drivers operate close to hardware, and thus implementing recovery routines may be particularly difficult in these subsystems. There may also be bugs in hardware, and existing observations indicate that many Linux device drivers tolerate such bugs poorly (Kadav09, ). The drivers subsystem has also been observed to be the most bug-prone one, including with respect to regression bugs (Chou01, ; Ruohonen24EASE, ). To that end, it might also be that hard assertions are used in the device driver subsystem to annotate code against verified and fixed bugs.
Finally, the dynamic AR(1) models too deserve a brief exposition. The CUSUM processes for these are visualized in Fig. 3. As can be seen, all fluctuation processes remain neatly within the boundaries. This observation reflects a fact that the differenced series are stationary; the means of the release-to-release change rates are constant. The coefficients for the lagged and differenced series are also modest in magnitude; with an exception of others, the coefficients are smaller than $0.2$ in absolute value. All in all, the dynamic AR(1) models signify that no major refactoring periods involving assertions have occurred. When backtracking to Fig. 1, the only exception is the sudden decline of hard assertions in the net subsystem late during the kernel’s third release series.
5. Conclusion
The paper examined the longitudinal evolution of assertions in the Linux kernel from its third release series to the ongoing sixth release series. Two research questions were postulated. The answers to these can be summarized as follows. With respect to RQ.1, the conclusion is that the kernel’s continuous growth, as being driven particularly by the addition of new device drivers, has been met with a growing use of softer assertion variants that produce warnings but do not cause an outright kernel panic. Regarding RQ.2, the conclusion is that at the same time the use of harder, panic-causing assertions has slowly decreased in some of the kernel’s major subsystems. The notable exceptions are the arch and drivers subsystems together with a leftover subsystem group labeled as others. In these subsystems the use of the harder BUG()-style assertions has remained stable or even increased. Nevertheless, it seems that a slow but visible trend is occurring in other Linux kernel subsystems toward handling errors in a more graceful way than causing a panic.
The handling of errors in the kernel space would also deserve more research. Indeed, whether and how well an operating system kernel handles bugs and fatal errors is a classical research question (Chou01, ). While there is a whole research branch on fault-tolerant operating systems and their kernels, it would be worthwhile to study also the error recovery strategies used in existing operating system kernels. This topic could be studied also empirically by focusing on frameworks and interfaces as well as programming patterns and idioms used for error handling in the kernel space. A comparative setup for empirical research could be constructed by again concentrating on the subsystems of the Linux kernel or enlarging the scope to other open source software operating system kernels. An interesting research path would open also by focusing on the potential differences between error handling with the C and Rust programming languages. Another option would be to continue with the themes noted in Section 2. For instance, the existing knowledge seems limited with respect to the alleged benefits of assertions for documenting code and helping debugging in the kernel context.

### v2

@@
 A Time Series Analysis of Assertions in the Linux Kernel
-Abstract.
+Abstract
 Assertions are a classical and typical software development technique.
 These are extensively used also in operating systems and their kernels, including the Linux kernel.
-The paper patches a gap in existing knowledge by empirically examining the longitudinal evolution of assertion use in the Linux kernel.
+The paper fills a gap in existing knowledge by empirically examining the longitudinal evolution of assertion use in the Linux kernel.
 According to the results, the use of assertions that cause a kernel panic has slightly but not substantially decreased from the kernel’s third to the sixth release series.
-At the same time, however, the use of softer assertion variants has increased; these do not cause a panic by default but instead produce warnings.
+At the same time, the use of softer assertion variants has increased; these do not cause a panic by default but instead produce warnings.
 With these time series results, the paper contributes to the existing but limited empirical knowledge base about operating system kernels and their long-term evolution.
-Design by contract; kernel panic; software evolution; subsystems
-ccs: Software and its engineering Operating systems
-1. Introduction
-Assertions are a classical programming technique used to validate a program’s execution at runtime.
-They also serve to document code for other developers.
-Assertions are typically used within functions to test preconditions placed at the beginning of a function and postconditions placed at the end of the function, although constraints can be placed also upon return values and intermediate states of the function (Rosenblum95, ).
+Keywords:
+D
+esign by contract, kernel panic, software evolution, subsystems
+1 Introduction
+Assertions are a classical programming technique used to verify a program’s execution at runtime.
+Their history traces all the way back to the early works by von Neumann and Turing [7].
+In addition to verification, assertions serve many other functions too.
+Among other things, they also serve to document code for other developers, which can be particularly valuable when operating with legacy systems and code [11].
+Assertions are typically used within functions to test preconditions placed at the beginning of a function and postconditions placed at the end of the function, although constraints can be placed also upon return values and intermediate states of the function [23].
+Assertions can be used also for testing many other conditions, such as whether a condition or a state is satisfied or reached at some time in the future [1, 7, 36].
 In software engineering and the computing world in general, assertions are also behind a classical debate between so-called defensive programming and design by contract.
-In general, the defensive variant seeks to particularly validate each and every input, whereas the latter assumes an implicit contract made explicit through assertions between a caller and a supplier.
-In terms of the design by contract paradigm, a supplier can assume that certain conditions are always guaranteed by a caller, whereas the caller can assume a safe output provided by the supplier.
-When operating with assertions instead of defensive programming practices, “any runtime violation of an assertion is not a special case but always the manifestation of a software bug”, tracing of which can be started by examining the given preconditions, implying “a bug in the client (caller)”, and the given postconditions associated with “a bug in the supplier (routine)” (Meyer92, , p.
-44).
-Regardless whichever side one takes in this debate, assertions are frequently used—and triggered—also in operating systems and their kernels.
+In general, the defensive variant does not trust anything coming from outside, thus seeking to particularly validate each and every input [31].
+In contrast, the design by contract programming philosophy assumes an implicit contract made explicit through assertions between a caller and a supplier; a supplier can assume that certain conditions are always guaranteed by a caller, whereas the caller can assume a safe output provided by the supplier.
+Some implementations have made the implicit contract visible by using a keyword assume for precondition checks and a keyword promise for postcondition checks [22].
+Given such checks, “any runtime violation of an assertion is not a special case but always the manifestation of a software bug”, tracing of which can be started by examining the given preconditions, implying “a bug in the client (caller)”, and the given postconditions associated with “a bug in the supplier (routine)” [17, p.
+44].
+Regardless whichever side one takes in this classical debate, assertions are frequently used—and triggered—also in operating systems and their kernels.
+The use is presumably explained partially by the context: given the overhead caused by defensive programming practices particularly for resource-constrained systems, assertions have the obvious benefit that these can be turned off during compilation [7].
 This background motives the paper’s focus on the Linux kernel.
-With respect to the Linux kernel, assertions have frequently been observed to be behind many of the crashes triggered by continuous kernel fuzzing frameworks (Ruohonen19RSDA, ).
+With respect to the Linux kernel, assertions have frequently been observed to be behind many of crashes triggered by kernel fuzzing frameworks [28].
 In the kernel context an explicitly triggered crash corresponds with a kernel panic.
 Without an assertion causing a panic, an alternative behavior might be that a whole operating system just hangs.
-Also these cases have historically been common in the Linux kernel (Gu03, ; Yoshimura12, ).
-Thus, the design by contract paradigm has a slightly different interpretation in the kernel context within which failing spectacularly may often, or at least sometimes, be a better option than to continue execution in undefined waters (FreeBSD24, ).
+Also these cases have historically been common in the Linux kernel [10, 37].
+Thus, the design by contract philosophy has a slightly different interpretation in the kernel context within which failing spectacularly may often, or at least sometimes, be a better option than to continue execution in undefined waters [2, 11].
 Different fallback and recovery solutions are often also difficult to design and implement in the kernel context.
 These and related points motivated a lively debate recently among kernel developers and other open source developers.
-The motivating argument put forward was that Linux kernel developers should slowly move away from using assertions that cause an outright panic; instead, they should prefer warnings and try to come up with means to recover from fatal errors (LWN24a, ).
+The motivating argument put forward was that Linux kernel developers should slowly move away from using assertions that cause an outright panic; instead, they should prefer warnings and try to come up with means to recover from fatal errors [8].
 In the Linux kernel the former panic-inducing assertions are used with different BUG()-style macros, whereas the latter are typically used with different WARN()-style variants.
 For a lack of a better terminology, these are hereafter known as “hard” and “soft” assertions.
-Regarding the hard assertions, the kernel documentation argues that “‘I’m too lazy to do error handling’ is not an excuse for using BUG()”, and that major “internal corruptions with no way of continuing may still use BUG(), but need good justification” (Kernel24a, ).
-With this guidance and the debate in mind, two research questions (RQs) are examined:
+Regarding the hard assertions, the kernel documentation argues that “‘I’m too lazy to do error handling’ is not an excuse for using BUG()”, and that major “internal corruptions with no way of continuing may still use BUG(), but need good justification” [33].
+With this guidance and the debate in mind, the following two research questions (RQs) are examined:
   • RQ.1: Has the kernel’s growing size been met with a growing use of soft assertions in some or all of its notable subsystems?
@@
 The final Section 5 concludes.
-2. Related Work
+2 Related Work
 In addition to the works already referenced, a couple of larger knowledge bases are worth mentioning.
@@
 Thus, regarding RQ.1, the software evolution research domain provides a good reference point.
-According to the so-called laws of software evolution, which were developed by Lehman and associates from the 1970s onward (Lehman97, ), many—if not the most—software artifacts are continuously changing, continuously growing, and continuously getting more complex (Ruohonen15JSEP, ).
-These laws correlate with the more recent concept of technical debt; it is necessary to pay back maintenance debt in order to continue to have a given software artifact in a reasonable state in the face of continuous growth and complexity.
-The Linux kernel has been seen as a prime example supporting the Lehman’s laws (Israeli10, ).
-Regarding growth, particularly new device drivers continuously added have been the driving force (Chou01, ; Ruohonen24EASE, ).
-Partially due to the growth, the Linux kernel is also continuously and heavily refactored, as manifested by the extensive amount of code churn (Ruohonen19RSDA, ).
+According to the so-called laws of software evolution, which were developed by Lehman and associates from the 1970s onward [16], many—if not the most—software artifacts are continuously changing, continuously growing, and continuously getting more complex [27].
+These laws correlate with the more recent concept of technical debt; it is necessary to pay back maintenance debt in order to continue to have a given software artifact in a reasonable state in the face of continuous growth and increasing complexity.
+The Linux kernel has been seen as a prime example supporting the Lehman’s laws [12].
+Regarding growth, particularly new device drivers continuously added have been the driving force [6, 25].
+Partially due to the growth, the Linux kernel is also continuously and heavily refactored, as manifested by the extensive amount of code churn [28].
 In other words, technical debt is also continuously paid back.
@@
 The second knowledge base is the somewhat limited amount of existing empirical research on assertions.
-This claim should not be interpreted to mean that empirical research would be absent; particularly the benefits of assertions in software testing have been investigated empirically (Taromirad23, ).
-In this regard, assertions have been observed to correlate with the effectiveness of test suites (Zhang15, ) and less buggy files (Kudrjavets06, ), among other things.
-Another empirical research area has focused on the benefits of assertions for troubleshooting and debugging, understandability, and documentation (Patel17, ; Peruma24, ).
-However, to the best of the author’s knowledge, no empirical studies have been conducted about the longitudinal evolution of the use of assertion in the Linux kernel or any other kernel for that matter.
-Thus, the paper fills a small but notable gap in existing knowledge.
-3. Research Setup
-3.1. Data and Time Series
+This claim should not be interpreted to mean that empirical research would be absent altogether; among other things, the benefits of assertions in software testing have been investigated empirically [5, 30].
+In this regard, assertions have been observed to correlate with the effectiveness of test suites [40] and less buggy files [15].
+Another empirical research area has focused on the benefits of assertions for troubleshooting and debugging, reusability, understandability, and documentation [18, 20, 21, 35].
+This research area aligns closely with the design by contract philosophy.
+However, to the best of the author’s knowledge, no empirical studies have been conducted about the longitudinal evolution of the use of assertion in the Linux kernel—or any other operating system kernel for that matter.
+Thus, the paper fills a small but notable gap in existing software engineering knowledge.
+3 Research Setup
+3.1 Data and Time Series
 The Linux kernel more or less follows the traditional semantic versioning scheme.
 Release series are the highest level of abstraction; currently, the Linux kernel operates with the sixth release series.
-Major releases, such as $6.1,6.2,6.3,\ldots$, are made based on a loosely time-based release process; a new major release occurs roughly every two or three months (Kernel24b, ).
+Major releases, such as $6.1,6.2,6.3,\ldots$, are made based on a loosely time-based release process; a new major release occurs roughly every two or three months [32].
 The adjective major signifies that such releases may contain everything from new features to internal changes in the kernel’s application programming interface.
@@
 Therefore, the major releases are more suitable for observing the evolution of assertions in the kernel.
-Thus, each major release in the third, fourth, fifth, and sixth release series was downloaded from the kernel’s online archives (Kernel24c, ).
+Thus, each major release in the third, fourth, fifth, and sixth release series was downloaded from the kernel’s online archives [34].
 At the time of writing (20 December 2024), the amounts of major releases released in these release series were 20, 21, 20, and 13, respectively.
@@
 Although the length is rather short in general, it is still sufficient for time series analysis with formal methodologies.
-Both the overall software architecture of the Linux kernel and its software development rely on different subsystems (Ruohonen24EASE, ).
+Both the overall software architecture of the Linux kernel and its software development rely on different subsystems [25].
 In what follows, the time series analysis is also based on major subsystems of the kernel.
 The major ones observed are: arch (architectural code dependent on instruction set architectures), drivers (device drivers), fs (file systems), net (networking), and sound (sound architectures).
-By following existing research (JiangJiang24, ; Ruohonen24EASE, ), these were proxied by the top-level directories in the kernel’s file system representation.
+By following existing research [13, 25], these were proxied by the top-level directories in the kernel’s file system representation.
 In addition, a category of others was used to merge other notable subsystems, such as kernel for general kernel code, crypto for cryptographic implementations, and security for the kernel’s security implementations.
@@
 The device driver subsystem is the largest one regardless of a size metric.
-Table 1. Average Sizes of the Subsystems Across Releases
+Table 1: Average Sizes of the Subsystems Across Releases
 Subsystem Files LOC
@@
 Then, for each subsystems, four time series were constructed:
-  1. (1)
+  1. 1.
     $\textmd{FILE}_{it}$ counts the number of files containing C programming code, as identified by a file’s ending to a suffix .c, in the $i$:th subsystem of the $t$:th major release.
-  2. (2)
+  2. 2.
     $\textmd{LOC}_{it}$ measures the lines of code, including code comments, of the files containing C code, as identified similarly to $\textmd{FILES}_{it}$, in the $i$:th subsystem of the $t$:th major release.
-  3. (3)
+  3. 3.
     $\textmd{BUG}_{it}$ measures the amount of hard assertions in the C files present in the $i$:th subsystem of the $t$:th major release.
 These were identified by parsing each file line by line and counting the occurrences of the character strings BUG( and BUG_ON(.
-  4. (4)
+  4. 4.
     $\textmd{WARN}_{it}$ measures the amount of soft assertions in the .c-suffixed files present in the $i$:th subsystem of the $t$:th major release.
 Analogously to $\textmd{BUG}_{it}$, these were identified by parsing each file line by line and counting the occurrences of the character strings WARN(, WARN_ON(, and WARN_ON_ONCE(.
-The computations are carried out with the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series as well scaled variants of the two time series.
+To give a short example about the use of the assertion variants in practice, the file ./drivers/base/regmap/regcache.c triggers a kernel panic in a function in case a precondition is not satisfied by a caller:
+        void regcache_exit(struct regmap *map)
+                if (map->cache_type == REGCACHE_NONE)
+                        return;
+                BUG_ON(!map->cache_ops);
+The above case is a good example because it could be argued that a softer assertion variant might work or that an explicit check could be done analogously to checking for the REGCACHE_NONE constant.
+The same file uses also BUG() as a default fallback case in a switch statement, indicating for a caller that only the values explicitly listed in the statement are assumed to be valid.
+Then, the computations are carried out with the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series as well scaled variants of the two time series.
 Two simple scaling functions are used: the first scales the plain series by the total amount of hard and soft assertions in the $i$:th subsystems of the $t$:th release, and the second by a subsystem’s size at a given major release.
 In other words:
-(1) $\overline{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{BUG}_{it}+\textmd{WARN}_{it}}\quad\textmd{and}\quad\widetilde{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{FILE}_{it}},$
+$\overline{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{BUG}_{it}+\textmd{WARN}_{it}}\quad\textmd{and}\quad\widetilde{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{FILE}_{it}},$ (1)
 and likewise for $\textmd{WARN}_{it}$.
 The scaling by $\textmd{FILE}_{it}$ is arbitrary; none of the results reported changed when scaled by $\textmd{LOC}_{it}$.
-Finally, it can be remarked that a replication package is available online (Ruohonen24rep, ).
-3.2. Methods
+Finally, it can be remarked that a replication package is available online [24].
+It contains the dataset and the code.
+3.2 Methods
 Statistical methods for cumulative sum (CUSUM) processes are suitable for examining the research questions.
-The history behind CUSUM processes is rich, starting from the 1950s and peaking in the 1970s during which first formal statistical tests were derived for testing structural changes in time series (Brown75, ).
+The history behind CUSUM processes is rich, starting from the 1950s and peaking in the 1970s during which first formal statistical tests were derived for testing structural changes in time series [3].
 By a structural change it is meant that a model estimated for a time series does not remain stable over time.
+To this end, a term parameter stability is sometimes used as an alternative [4].
 The concept is much broader than the concept of (covariance) stationary, which only requires that a mean and a variance of a time series remain constant.
 However, in the present context these two concepts are closely related because the focus is on the means of $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$.
-With respect to RQ.2, a weak requirement is that the mean of the former should not remain constant.
-A strong requirement implies that the means of both series should not remain constant because developers are encouraged to move from hard assertions to the softer variants.
+With respect to RQ.2, a weak requirement for a positive answer is that the means of the former should not remain constant for all or at least most $i$.
+A strong requirement implies that the means of both series should not remain constant in all or most subsystems because developers are encouraged to move from hard assertions to the softer variants.
 The CUSUM processes are thus computed for both “static mean models” and “dynamic autoregression models”, both estimated with standard ordinary least squares (OLS).
@@
 The underlying (null) hypotheses are that $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{74}$ and $\sigma^{2}_{\epsilon_{1}}=\sigma^{2}_{\epsilon_{2}}=\cdots=\sigma^{2}_{\epsilon_{74}}$, and likewise for $\beta_{t}$ and $\sigma^{2}_{\varepsilon_{t}}$.
-In particular, the conditional means should remain constant, although the word conditional is a slight misnomer in the present context because only constants are included in the models.
+In particular, the conditional means should remain constant, although the word conditional is a slight misnomer in the present context because only constants are included.
 The dynamic first-order autoregression, denoted by AR(1), models are estimated for differenced series, $\Delta\textmd{BUG}_{it}=\textmd{BUG}_{it}-\textmd{BUG}_{it-1}$ and $\Delta\textmd{WARN}_{it}$, given that many of the series are not stationary.
@@
 The CUSUM processes are particularly useful as visual diagnostic tools.
-The strucchange package (Zeileis02, ) for R is used for the visualizations and associated computations.
+The strucchange package [39] for R is used for the visualizations and associated computations.
 It allows to visualize the CUSUM processes computed with visual boundaries; 99% confidence level boundaries are used for the visualizations presented.
-In addition, the package offers more formal so-called M-fluctuation tests (Zeileis06, ).
+With the earlier points in mind and using the eventually assertion notation proposed in the literature [1], the weak requirement for a positive answer to RQ.2 means that eventually("the boundaries are crossed" $\equiv$ TRUE) for all or at least most $i$.
+The logic is thus similar to using CUSUM processes for anomaly detection [19].
+In addition, the strucchange package offers more formal so-called M-fluctuation tests [38].
 These are used to compute analogous models than the static (OLS) mean models used for the visualizations except that a Poisson regression is used for the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series.
@@
 In addition, to examine RQ.1 in a little more detail, a simple non-linear growth curve estimation is carried out.
-Although numerous growth curves have been specified in the literature (Ruohonen15COSE, ), the standard logistic growth curve is sufficient for the present purposes.
-It is written the growthcurver package (Sprouffske16, ) used for computation as:
-(2) $y_{t}=\frac{\alpha}{1+\frac{\alpha-y_{0}}{y_{0}}e^{-\beta t}},$
+Although numerous growth curves have been specified in the literature [26], the standard logistic growth curve is sufficient for the present purposes.
+It is written in the growthcurver package [29] used for computation as:
+$y_{t}=\frac{\alpha}{1+\frac{\alpha-y_{0}}{y_{0}}e^{-\beta t}},$ (2)
 where $y_{t}$ denotes a time series estimated with non-linear least squares, $\alpha$ is the asymptotic maximum, $y_{0}$ is the initial value for a growth, and $\beta$ is the growth rate.
@@
 As the logistic growth curve is ill-suited for such series, the growth rate coefficients are interpreted together with their standard errors.
-4. Results
+4 Results
 The presentation of the results can be started by taking a look at the time series across the six subsystems.
@@
 Although a reason can be left for further examinations, the observation could be interpreted in a positive light in a sense that machine-dependent code has slightly decreased.
-Figure 1. $\textmd{FILE}_{it}$, $\textmd{BUG}_{it}$, and $\textmd{WARN}_{it}$
-Table 2. Logistic Growth Rate Parameters and Their Standard Errors (SEs)1
+Figure 1: $\textmd{FILE}_{it}$, $\textmd{BUG}_{it}$, and $\textmd{WARN}_{it}$
+Table 2: Logistic Growth Rate Parameters and Their Standard Errors (SEs)1
 $\textmd{FILE}_{it}$ $\textmd{LOC}_{it}$ $\textmd{BUG}_{it}$ $\textmd{WARN}_{it}$
 $\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$) $\hat{\beta}$ SE($\hat{\beta}$)
-Subsystem arch $<$0.0001 $<$0.0001 0.0534 0.0015
+Subsystem arch <<0.0001 <<0.0001 0.0534 0.0015
 drivers 0.0739 0.0028 0.0648 0.0026 0.0676 0.0027
-fs 0.0306 0.0033 0.0413 0.0028 $<$0.0001 $<$0.0001 0.0553 0.0033
-net 0.0781 0.0044 0.0641 0.0024 $<$0.0001 $<$0.0001 0.0606 0.0020
+fs 0.0306 0.0033 0.0413 0.0028 <<0.0001 <<0.0001 0.0553 0.0033
+net 0.0781 0.0044 0.0641 0.0024 <<0.0001 <<0.0001 0.0606 0.0020
 sound 0.0574 0.0028 0.0594 0.0022 0.2722 0.0336
@@
   • 1 Estimates are not shown in case a standard error is larger than a parameter estimate.
-Table 3. Generalized M-fluctuation Tests ($p$-values)
+Table 3: Generalized M-fluctuation Tests ($p$-values)
 Series $\textmd{BUG}_{it}$ $\overline{\textmd{BUG}}_{it}$ $\widetilde{\textmd{BUG}}_{it}$ $\textmd{WARN}_{it}$ $\overline{\textmd{WARN}}_{it}$ $\widetilde{\textmd{WARN}}_{it}$
 Model Poisson OLS OLS Poisson OLS OLS
-Subsystem arch $0.0003$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
-drivers $0.0002$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
-fs $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
-net $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
-sound $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
-others $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $0.0001$
-Figure 2. OLS-CUSUM Processes for $\textmd{FILE}_{it}$ and $\textmd{BUG}_{it}$ (the circled boundaries denote a $p<0.001$ confidence level)
-Figure 3. OLS-CUSUM Processes for AR(1) models of $\Delta\textmd{FILE}_{it}$ and $\Delta\textmd{BUG}_{it}$ (the boundaries denote a $p<0.001$ confidence level)
+Subsystem arch $0.0003$ <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
+drivers $0.0002$ <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
+fs <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
+net <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
+sound <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
+others <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 $0.0001$
+Figure 2: OLS-CUSUM Processes for $\textmd{FILE}_{it}$ and $\textmd{BUG}_{it}$ (the circled boundaries denote a $p<0.001$ confidence level)
+Figure 3: OLS-CUSUM Processes for AR(1) models of $\Delta\textmd{FILE}_{it}$ and $\Delta\textmd{BUG}_{it}$ (the boundaries denote a $p<0.001$ confidence level)
 With these observations about growing sizes of most subsystems, the logistic growth curve estimates can be subsequently considered.
@@
 In general, the answer to RQ.1 is positive; in most major subsystems the growing sizes have been met with equally growing uses of soft assertions.
-Having said that, it remains a subject of a debate whether the growth rates of $\textmd{WARN}_{it}$ should be even higher in order to satisfy the ever-growing sizes of some subsystems, the implicit design by contract principle, and the overall goal of providing non-fatal debugging information to users.
+Having said that, it remains a subject of a debate whether the growth rates of $\textmd{WARN}_{it}$ should be even higher in order to satisfy the ever-growing sizes of some subsystems, the implicit design by contract principle, and the overall goal of providing non-fatal debugging information.
 Turning to the CUSUM processes, the formal M-fluctuation tests are summarized in Table 3.
@@
 A reasonable presumption might be that both arch and drivers operate close to hardware, and thus implementing recovery routines may be particularly difficult in these subsystems.
-There may also be bugs in hardware, and existing observations indicate that many Linux device drivers tolerate such bugs poorly (Kadav09, ).
-The drivers subsystem has also been observed to be the most bug-prone one, including with respect to regression bugs (Chou01, ; Ruohonen24EASE, ).
+There may also be bugs in hardware, and existing observations indicate that many Linux device drivers tolerate such bugs poorly [14].
+The drivers subsystem has also been observed to be the most bug-prone one, including with respect to regression bugs [6, 25].
 To that end, it might also be that hard assertions are used in the device driver subsystem to annotate code against verified and fixed bugs.
@@
 1, the only exception is the sudden decline of hard assertions in the net subsystem late during the kernel’s third release series.
-5. Conclusion
+5 Conclusion
 The paper examined the longitudinal evolution of assertions in the Linux kernel from its third release series to the ongoing sixth release series.
@@
 The handling of errors in the kernel space would also deserve more research.
-Indeed, whether and how well an operating system kernel handles bugs and fatal errors is a classical research question (Chou01, ).
+Indeed, whether, how, and how well an operating system kernel handles bugs and fatal errors are classical research questions [6, 9].
 While there is a whole research branch on fault-tolerant operating systems and their kernels, it would be worthwhile to study also the error recovery strategies used in existing operating system kernels.
-This topic could be studied also empirically by focusing on frameworks and interfaces as well as programming patterns and idioms used for error handling in the kernel space.
+The topic could be studied also empirically by focusing on frameworks and interfaces as well as programming patterns and idioms used for error handling in the kernel space.
 A comparative setup for empirical research could be constructed by again concentrating on the subsystems of the Linux kernel or enlarging the scope to other open source software operating system kernels.
@@
 For instance, the existing knowledge seems limited with respect to the alleged benefits of assertions for documenting code and helping debugging in the kernel context.
+Answers to these and related questions might help at reaching the goal of moving toward software assertion use in the Linux kernel.